# Small-scale magnetohydrodynamic dynamos: from deterministic chaos to turbulence

### v2

@@ I Introduction
$E(k)\propto\exp-(k/k_{\beta})^{\beta}.$ (1)1( 1 )
@@ II.1 Magnetic helicity
$\hat{h}=h_{m}+2{\bf B_{0}}\cdot\langle{\bf A}\rangle$ (4)4( 4 )
@@ II.3 Spontaneous breaking of local reflectional symmetry
${\rm I^{\pm}}=\lim_{V\rightarrow\infty}\frac{1}{V}\sum_{j}H_{j}^{\pm}$ (11)11( 11 )
@@ II.4 Examples
$\partial_{t}{\bf u}+({\bf u}\cdot\nabla){\bf u}=-\nabla\left(p+\frac{{\bf b}^{2}}{2}\right)+\left({\bf b}\cdot\nabla\right){\bf b}+\nu\nabla^{2}{\bf u}+{\bf f}$ (13)13( 13 )
$\partial_{t}{\bf b}=-\left({\bf u}\cdot{\bf\nabla}\right){\bf b}+\left({\bf b}\cdot\nabla\right){\bf u}+\eta\nabla^{2}\bf{b}$ ($)14$
$\nabla\cdot{\bf u}=0,~{}~{}~{}{\bf\nabla}\cdot{\bf b}=0,$ (15)15( 15 )
were solved in a periodic spatial cube.
@@ III Magneto-inertial range of scales
$B_{c}\propto\varepsilon_{h}^{1/2}~{}\varepsilon^{1/6}~{}k_{c}^{1/6}$ (16)16( 16 )
@@
$E(k)\propto\int_{0}^{\infty}P(k_{c})\exp-(k/k_{c})dk_{c}\propto\exp-(k/k_{\beta})^{\beta}$ (18)18( 18 )
@@
$P(k_{c})\propto k_{c}^{-1+\beta/[2(1-\beta)]}~{}\exp(-\gamma k_{c}^{\beta/(1-\beta)})$ (19)19( 19 )
@@
$\beta=\frac{2\alpha}{1+2\alpha}$ (20)20( 20 )
@@ IV.1 Dissipative MHD invariants
$\frac{\partial\langle z_{L}^{\pm}z_{L}^{\pm^{\prime}}\rangle}{\partial t}=\left(\frac{\partial}{\partial r}+\frac{4}{r}\right)C^{\pm}_{LLL}(r)+2\left(\frac{\partial^{2}}{\partial r^{2}}+\frac{4}{r}\right)D_{LL}$ (22)22( 22 )
@@
$C^{\pm}_{LLL}(R)\left.~{}~{}~{}~{}\rm{and}~{}~{}~{}~{}\frac{\partial D_{LL}}{\partial r}\right\arrowvert_{r=R}\rightarrow 0$
@@ V Hall magnetohydrodynamics
$\partial_{t}{\bf u}+({\bf u}\cdot\nabla){\bf u}=-\nabla\left(p+\frac{{\bf b}^{2}}{2}\right)+\left({\bf b}\cdot\nabla\right){\bf b}+\nu\nabla^{2}{\bf u}+{\bf f}$ (34)34( 34 )
$\partial_{t}{\bf b}=-\left({\bf u}\cdot{\bf\nabla}\right){\bf b}+\left({\bf b}\cdot\nabla\right){\bf u}+\eta\nabla^{2}\bf{b}+$
$+d_{i}\left({\bf j}\cdot{\bf\nabla}\right){\bf b}-d_{i}\left({\bf b}\cdot\nabla\right){\bf j}$ (35)35( 35 )
$\nabla\cdot{\bf u}=0,~{}~{}~{}{\bf\nabla}\cdot{\bf b}=0,$ (36)36( 36 )
@@ VI Geodynamo
$\partial_{t}{\bf u}+\left(\frac{2}{E}\,{\bf e_{z}}+\nabla\times{\bf u}\right)\times{\bf u}=-{\bf\nabla}p+\Delta{\bf u}+$
+(×$+(\nabla\times{\bf b})\times{\bf b}-\frac{Ra}{\beta}~{}C~{}\frac{D}{r_{o}}{\bf r}$ (37)37( 37 )
$\partial_{t}{\bf b}=\nabla\times({\bf u}\times{\bf b})+\frac{1}{Pr_{m}}\Delta{\bf b}$ (38)38( 38 )
$\partial_{t}C+{\bf u}\cdot{\bf\nabla}(C+C_{0})=\frac{1}{Pr}\Delta C$ (39)39( 39 )
$\nabla\cdot{\bf u}=0,~{}~{}~{}{\bf\nabla}\cdot{\bf b}=0$ (40)40( 40 )
@@
Figure 11 shows the geomagnetic power spectrum calculated using local subgrids from the World Magnetic Anomaly Map of the National Geophysics Data Center (WMAM) maus1 ). The data were detrended and azimuthally averaged, and the mean field was subtracted by the author of the Ref. maus2 (the spectral data were taken from Fig. 4 of the Ref. maus2 ). The land and ocean averages were considered separately (here $R$ is the Earth’s radius). The dashed curve indicates the helical spectral law Eq. (9) for the ocean data, whereas at the land the spectral law Eq. (21) indicates the magneto-inertial range of scales.
@@
In the paper Ref. maus1 a large-scale globally averaged total intensity spectrum computed using the spherical harmonic coefficients of the magnetic potential from a C89 global geomagnetic field model C89 cain was also presented. It is interesting to compare the spectral data with the above-suggested spectra. Figure 12 shows the spectral data taken from Fig. 2 of the Ref. maus1 . The dashed curves indicate the helical spectrum Eq. (9) (here $R$ is the Earth’s radius).
The C89 model used the observational data available at the time of its creation. Therefore it is useful to look at the modern geomagnetic data including, in particular, the low Earth orbit space-based observations. An example of such a global geomagnetic energy spectrum is given in Fig. 13 (in the semi-log scales). The spectral data were taken from Fig. 13 (the main field) of a recent paper Ref. styp . The dashed curves indicate the helical spectrum Eq. (9) (here $R$ is the Earth’s radius).
@@ VII.1 Numerical simulations
Figure 13: Global geomagnetic energy spectrum at the Earth’s surface in 2014.0.
@@
Figure 14 shows the magnetic energy spectra for a photosphere layer at different times of the dynamo evolution. The spectral data were taken from Fig. 3b of the Ref. kita ($k_{h}$ corresponds to the horizontal wavenumber). The two earlier time spectra ($t=1$h and $t=2$h in the simulation terms) are well-fitted by the exponential spectrum Eq. (2) ($\beta=1$, deterministic chaos). Next spectrum ($t=3$h) can be fitted by the Eq. (33) ($\beta=5/6$, dissipative distributed chaos dominated by the magnetic Loitsianskii invariant), then the spectrum at $t=4$h can be fitted by Eq. (31) ($\beta=3/4$, dissipative distributed chaos dominated by the magnetic Birkhoff-Saffman invariant).
@@ VII.3 The role of a mean magnetic field
$B_{c}\propto\varepsilon_{\hat{h}}^{1/2}~{}(\varepsilon\widetilde{B}_{0})^{1/8}k_{c}^{1/8}$ (42)42( 42 )

### v3

@@ I Introduction
$E(k)\propto\exp-(k/k_{\beta})^{\beta}.$ (1)
@@ II.1 Magnetic helicity
$\hat{h}=h_{m}+2{\bf B_{0}}\cdot\langle{\bf A}\rangle$ (4)
@@ II.3 Spontaneous breaking of local reflectional symmetry
${\rm I^{\pm}}=\lim_{V\rightarrow\infty}\frac{1}{V}\sum_{j}H_{j}^{\pm}$ (11)
@@ II.4 Examples
$\partial_{t}{\bf u}+({\bf u}\cdot\nabla){\bf u}\!\!$ $=$ $\!\!-\nabla P+\left({\bf b}\cdot\nabla\right){\bf b}+\nu\nabla^{2}{\bf u}+{\bf f}$ (13)
$\partial_{t}{\bf b}+\left({\bf u}\cdot{\bf\nabla}\right){\bf b}\!\!$ $=$ $\!\!\left({\bf b}\cdot\nabla\right){\bf u}+\eta\nabla^{2}\bf{b}$ (14)
$\nabla\cdot{\bf u}$ $=$ $0,~{}~{}~{}{\bf\nabla}\cdot{\bf b}=0,$ (15)
were solved in a periodic spatial cube (here $P=p+\frac{{\bf b}^{2}}{2}$ and $p$ is the kinetic pressure normalized by the constant density).
@@ III Magneto-inertial range of scales
$B_{c}\propto\varepsilon_{h}^{1/2}~{}\varepsilon^{-1/6}~{}k_{c}^{1/6}$ (16)
@@
$E(k)\propto\int_{0}^{\infty}P(k_{c})\exp-(k/k_{c})dk_{c}\propto\exp-(k/k_{\beta})^{\beta}$ (18)
@@
$P(k_{c})\propto k_{c}^{-1+\beta/[2(1-\beta)]}~{}\exp(-\gamma k_{c}^{\beta/(1-\beta)})$ (19)
@@
$\beta=\frac{2\alpha}{1+2\alpha}$ (20)
@@ IV.1 Dissipative MHD invariants
$\frac{\partial\langle z_{L}^{\pm}z_{L}^{\pm^{\prime}}\rangle}{\partial t}=\left(\frac{\partial}{\partial r}+\frac{4}{r}\right)C^{\pm}_{LLL}(r)+2\left(\frac{\partial^{2}}{\partial r^{2}}+\frac{4}{r}\right)D_{LL}$ (22)
@@
$C^{\pm}_{LLL}(R)\left.~{}~{}~{}~{}\rm{and}~{}~{}~{}~{}\frac{\partial D_{LL}}{\partial r}\right\arrowvert_{r=R}\rightarrow 0$ (24)
@@ V Hall magnetohydrodynamics
$\partial_{t}{\bf u}+({\bf u}\cdot\nabla){\bf u}\!\!$ $=$ $\!\!-\nabla P+\left({\bf b}\cdot\nabla\right){\bf b}+\nu\nabla^{2}{\bf u}+{\bf f}$ (35)
$\partial_{t}{\bf b}+\left({\bf u}\cdot{\bf\nabla}\right){\bf b}\!\!$ $=$ $\!\!\left({\bf b}\cdot\nabla\right){\bf u}+\eta\nabla^{2}\bf{b}$ (36)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!+d_{i}\left({\bf j}\cdot{\bf\nabla}\right){\bf b}-d_{i}\left({\bf b}\cdot\nabla\right){\bf j}$
$\nabla\cdot{\bf u}$ $=$ $0\displaystyle 0,~{}~{}{\bf\nabla}\cdot{\bf b}=$ 0 (37)
@@ VI Geodynamo
$\partial_{t}{\bf u}+({\bf u}\cdot\nabla){\bf u}$ $=$ $-{\bf\nabla}p-\left(\frac{2}{E}\,{\bf e_{z}}\times{\bf u}\right)+\Delta{\bf u}+$ (38)
+(×$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!+(\nabla\times{\bf b})\times{\bf b}-\frac{Ra}{\beta}~{}C~{}\frac{D}{r_{o}}{\bf r}$
$\partial_{t}{\bf b}+\left({\bf u}\cdot{\bf\nabla}\right){\bf b}$ $=$ $\left({\bf b}\cdot\nabla\right){\bf u}+\frac{1}{Pr_{m}}\Delta{\bf b}$ (39)
$\partial_{t}C+{\bf u}\cdot{\bf\nabla}(C+C_{0})$ $=$ $\frac{1}{Pr}\Delta C$ (40)
$\nabla\cdot{\bf u}=0,~{}~{}~{}{\bf\nabla}\cdot{\bf b}$ $=$ 00 0 (41)
@@
Another example of a numerical simulation trying to approach the real conditions can be found in a recent paper Ref. sffj . In this numerical simulation, the modeled liquid was enclosed in a rotating spherical shell between two radii $r_{o}$ and $r_{i}$ and the ratio $r_{i}/r_{o}=0.35$ (as for the Earth, cf the example of numerical simulation in Introduction). Both boundaries of the shell were taken no-slip and impermeable. The outer boundary was electrically insulating, and the inner solid core had the same conductivity as the liquid outer core. The inner core temperature was kept constant. Whereas the gradient of the temperature was kept constant and a uniform heat source was distributed throughout the outer core. The constant heat flux boundary conditions were applied at the core-mantle boundary. The dynamo was initiated and supported by the thermal convection under the strong influence of the rotation (the Ekman number $E=\nu/\Omega(r_{o}-r_{i})^{2}=3\times 10^{-7}$, the magnetic Prandtl number $Pr_{m}=\nu/\eta=0.05$).
Figure 11 shows the magnetic energy spectrum obtained at this numerical simulation (the spectral data were taken from Fig. 3b of the Ref. sffj . The spectrum was time and $r$ averaged over the whole shell, $R$ is an effective radius ($r_{i}<R<r_{o}$) and the wave number $k=\sqrt{l(l+1)}/R$ where $l$ is the spherical harmonic degree.
The dashed curve indicates the spectral law Eq. (21), i.e. the magneto-inertial range for the small scales. One can see that the small-scale dynamo coexists with the large-scale one in this simulation as well.
Figure 12 shows the geomagnetic power spectrum calculated using local subgrids from the World Magnetic Anomaly Map of the National Geophysics Data Center (WMAM) maus1 ). The data were detrended and azimuthally averaged, and the mean field was subtracted by the author of the Ref. maus2 (the spectral data were taken from Fig. 4 of the Ref. maus2 ). The land and ocean averages were considered separately (here $R$ is the Earth’s mean radius). The dashed curve indicates the helical spectral law Eq. (9) for the ocean data, whereas at the land the spectral law Eq. (21) indicates the magneto-inertial range of scales.
@@
Figure 11: Time-averaged magnetic energy spectrum at the Ekman number $E=3\times 10^{-7}$ and $Pr_{m}=0.05$ (numerical simulation).
@@
Figure 14: Global geomagnetic energy spectrum at the Earth’s surface in the period 1999-2020.
In the paper Ref. maus2 a large-scale globally averaged total intensity spectrum computed using the spherical harmonic coefficients of the magnetic potential from a C89 global geomagnetic field model C89 cain was also presented. It is interesting to compare the spectral data with the above-suggested spectra. Figure 13 shows the spectral data taken from Fig. 2 of the Ref. maus2 . The dashed curves indicate the helical spectrum Eq. (9) (here $R$ is the Earth’s radius).
Figure 14 shows the spectral data taken from Fig. 2 of a recent paper finlay ). The spectral computations were produced using the CHAOS-7 model of the near-Earth magnetic field for the period 1999-2020 based on the observations made by the low-Earth orbit satellites CryoSat-2, Swarm, CHAMP, SAC-C, Orsted, and SAC-C, and on the ground observatory measurements. The dashed curve indicates the helical spectrum Eq. (9) (here $R$ is the Earth’s mean radius).
@@ VII.1 Numerical simulations
Figure 15 shows the magnetic energy spectra for a photosphere layer at different times of the dynamo evolution. The spectral data were taken from Fig. 3b of the Ref. kita ($k_{h}$ corresponds to the horizontal wavenumber). The two earlier time spectra ($t=1$h and $t=2$h in the simulation terms) are well-fitted by the exponential spectrum Eq. (2) ($\beta=1$, deterministic chaos). Next spectrum ($t=4$h) can be fitted by the Eq. (33) ($\beta=5/6$, dissipative distributed chaos dominated by the magnetic Loitsianskii invariant), then the spectrum at $t=5$h can be fitted by Eq. (31) ($\beta=3/4$, dissipative distributed chaos dominated by the magnetic Birkhoff-Saffman invariant).
@@ VII.3 The role of a mean magnetic field
$B_{c}\propto\varepsilon_{\hat{h}}^{1/2}~{}(\varepsilon\widetilde{B}_{0})^{-1/8}k_{c}^{1/8}$ (43)